OpenStudy (anonymous):

@tHe_FiZiCx99 what is -4 1/2 times -6 2/3

OpenStudy (the_fizicx99):

I got 18 2/3

OpenStudy (anonymous):

-24 1/3 24 24 1/3 -30

OpenStudy (anonymous):

thoes r the answers i really need help

OpenStudy (solomonzelman):

\[-4\frac{1}{2}-6\frac{2}{3}\]using \[a\frac{b}{c}=\frac{(a \times c) + b}{c},\]\[-\frac{9}{2}-\frac{15}{3}\]\[-\frac{27}{6}-\frac{30}{6}\]\[-\frac{57}{6}\]\[-24\frac{1}{2}\]THE choices are wrong.

OpenStudy (anonymous):

what is -3 1/3 divided by 9??